# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v1

@@ IV.2 Boson and fermion momentum distributions
For generic coupling strength, the numerical Fermi momentum distributions are shown in Fig. 12 at different bosonic concentrations $x$. We only report the case $\eta_{\mathrm{B}}=0$, since the effect of the BB repulsion is only limited to a slight renormalization of the Fermi step. We find a qualitatively analogous phenomenology to Fig. 11 even though the latter is obtained in the strong coupling limit of the theory. This is because the effective interaction acting on fermions is given by an interplay between $g$ and $x$, that is, even if $g$ is strong, reducing $x$ results in an effectively weaker $g$. In addition, we observe that the position of the Fermi step does not change significantly from weak to strong coupling, in contrast to the 3D case where the Fermi sphere is emptied by the vanishing of the Fermi wave vector [83]. In 2D, the destruction of the Fermi surface occurring with coupling or concentration takes place due to a strong renormalization of the fermionic quasi-particles, as described in Sec. III, and it is their weight rather than their associated Fermi momentum that vanishes.
@@
IV.3 Boson and fermion chemical potentials
@@
Figure 13: (a): Bosonic chemical potential $\mu_{\mathrm{B}}$ (shifted by the binding energy ${\varepsilon_{0}}$) in units of the Fermi energy $E_{\mathrm{F}}$ as a function of the BF coupling $g$, for different values of bosonic concentration $x$ at zero boson-boson repulsion $\eta_{\mathrm{B}}=0$. Dotted lines on the left: weak-coupling expression (58). Dotted lines on the right: strong coupling asymptotics of Sec. III. Inset: comparison between $-\mu_{\mathrm{B}}/{E_{\mathrm{F}}}$ and the binding energy ${\varepsilon_{0}}/{E_{\mathrm{F}}}$ (dotted line). (b): Corresponding fermionic chemical potential $\mu_{\mathrm{F}}$ in units of the Fermi energy $E_{\mathrm{F}}$. Dotted lines on the left: weak-coupling expression (59) (reported for clarity only for the two extreme values of $x$). Dotted lines on the right: strong coupling asymptotics of Sec. III.
@@
The first term in Eq. (58) arises from the standard Bogoliubov theory in 2D [91] whilst the second contribution stems from the BF interaction. One sees in Fig. 13 that in the weak-coupling regime, the numerical curves for both chemical potentials neatly approach the asymptotic limits (58,59) (dotted line on the left side). For the bosonic chemical potential this occurs in a universal way, i.e., independently of the value of $x$, as required by Eq. (58) which does not depend on the bosonic concentration (and ${\varepsilon_{0}}$ is exponentially suppressed in this regime). In the opposite regime, both chemical potentials approach their strong-coupling benchmarks, depicted by the dotted lines on the right side of the graph, which are seen to be strongly dependent on concentration. Note that, for the boson chemical potential, Fig. 13(a) reports the quantity ${\mu_{\mathrm{B}}}+{\varepsilon_{0}}$, to subtract from the bosonic chemical potential the leading strong-coupling contribution, which is given by $-{\varepsilon_{0}}$, while the inset avoids this subtraction and illustrates on a logarithmic scale how the bosonic chemical potential tends to the binding energy, thus showing that all bosons are paired at large values of $g$.
Summing up, we find that two interaction regimes emerge from the above analysis: one for small $g$ where barely dressed fermionic atoms are found along with weakly interacting bosons and varying the concentration has little effect, and one of large $g$ in which, as $x$ increases, the atomic assembly is strongly renormalized and experiences a collapse of the Fermi surface. The boundary between them can be roughly estimated by looking at the smallest coupling at which ${\mu_{\mathrm{F}}}$ vanishes as a function of $x$ and turns out to correspond to about $g\simeq 1$.
@@
Finally, we note that we have verified that at finite $\eta_{\mathrm{B}}=0.1$ the results remain qualitatively the same, with quantitative changes affecting only ${\mu_{\mathrm{B}}}$, as also expected from the weak-coupling expression Eq. (58). The fermionic chemical potential is instead unaffected, in line with Eq. (59). On the other hand, in the strong coupling regime the boson-boson repulsion is essentially irrelevant due to the dominant dimer-atom repulsion.
@@
Figure 14: (a): Bosonic condensate fraction $n_{0}/n_{\mathrm{B}}$ as a function of the BF coupling parameter $g$, for different values of boson concentration $x$ at zero BB repulsion $\eta_{\mathrm{B}}=0$. Circles: Diagrammatic Monte Carlo results for the polaron quasiparticle residue $Z$ [105]. Squares: non-self-consistent $T$-matrix approximation results for the polaron quasiparticle residue $Z$ [106]. (b): Square of hybridization energy scale $\Delta_{0}$ (in units of $E_{\mathrm{F}}^{2}$), as a function of the BF coupling $g$, for different values of the bosonic concentration $x$ and boson-boson repulsion $\eta_{\mathrm{B}}=0$. Strong coupling benchmarks from Sec. III are also reported as dotted lines on the right.
@@
IV.4 Boson condensate density and atom-molecule hybridization energy scale
@@
Figure 14(a) also reports the results for the polaron quasiparticle residue $Z$ previously obtained with Diagrammatic Monte Carlo simulations (circles) [105] and within a non-self-consistent $T$-matrix approach [106] (squares). This is because in 3D it was found [88, 33] that the universal condensate fraction essentially matches the quasiparticle residue $Z$ of the Fermi polaron (which in our Bose-Fermi mixture corresponds to the limit of a single boson immersed in a Fermi gas).
@@
Such identification between the universal condensate fraction and $Z$ is challenged by the present results in 2D. One sees that, even though the condensate fraction $n_{0}/n_{\rm B}$ and the polaron quasiparticle residue $Z$ share the same qualitative behavior, significant quantitative discrepancies between the two quantities are present. We emphasize that the existence of such a discrepancy cannot be ascribed to the specific choice of the self-energy used in our calculation. One sees indeed that such a discrepancy occurs also when $Z$ is obtained with a non-self-consistent $T$-matrix approach (squares in Fig. 14(a)), which is exactly the same approach used here when carried over to the single-impurity limit. We thus conclude that the equivalence between the quasiparticle residue $Z$ of the Fermi polaron and the universal condensate fraction observed in 3D is only approximate, and the differences between these two quantities are amplified in 2D.
A further difference with respect to the 3D case is observed in the behavior of $n_{0}$ for strong coupling. While in 3D $n_{0}$ drops identically to zero for values of $g$ larger than a critical coupling $g_{c}$, in 2D $n_{0}$ never vanishes (even though it becomes exponentially small with the coupling $g$). This is because, as discussed in Sec. III.4, the energy scale $\Delta_{0}\propto(n_{0}{\varepsilon_{0}})^{1/2}$ has a non-negligible effect even when $n_{0}/{n_{\mathrm{B}}}\approx 0$, if the boson and fermion populations are not exactly balanced. One sees indeed in Fig. 14(b) that $\Delta_{0}$ vanishes only in the limit of balanced populations $x=1$. Figure 14(b) also shows that the hybridization energy scale $\Delta_{0}$ is almost negligible for small $g$ whereas, for large $g$, $\Delta_{0}$ saturates at values that depend on $x$. In this latter case, the maximum value is found at about $x\simeq 0.7$ (short-dashed line) in line with the change of sign of ${\mu_{\mathrm{F}}}$ in Fig. 13 and Fig. 3.
IV.5 Tan’s contact parameter and composite-fermion density
@@
For large momenta the fermionic and bosonic distribution functions have the following asymptotic behaviors
${n_{\mathrm{F}}}\left(k\gg k_{\mathrm{F}}\right)=\frac{{{C_{\mathrm{BF}}}/4m_{r}^{2}}}{\left(\frac{k^{2}}{2m_{r}}-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}\right)^{2}},$ (60)
${n_{\mathrm{B}}}(k\gg{k_{\mathrm{F}}})=\frac{{{C_{\mathrm{BF}}}/4m_{r}^{2}}}{\left(\frac{k^{2}}{2m_{r}}-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}\right)^{2}}+\frac{\Sigma^{2}_{12}}{4\left(\frac{k^{2}}{2m_{\mathrm{B}}}-{\mu_{\mathrm{B}}}\right)^{2}},$ (61)
@@
whose leading order term, proportional to $k^{-4}$ defines the Tan’s contact parameter $C_{s}=n_{s}(k)k^{4},s=\mathrm{F,B}$, for fermions and bosons respectively [100, 101, 102]. The coefficient ${C_{\mathrm{BF}}}$ gives the contribution to the Tan’s contacts due to BF pairing. Within our $T$-matrix self-energy approach, it can be shown to be given by
${C_{\mathrm{BF}}}$ $=4m_{r}^{2}\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\int\frac{d\Omega}{2\pi}T({\mathbf{P}},\Omega)e^{i\Omega 0^{+}}$ (62)
@@
Figure 15: (a): Dimensionless Tan’s contact constant ${C_{\mathrm{BF}}}/m_{\mathrm{F}}^{2}E_{\mathrm{F}}^{2}$ as a function of the BF coupling $g$, for different values of the bosonic concentration $x$ at boson-boson repulsion $\eta_{\mathrm{B}}=0$. Inset: same data divided by the boson concentration $x$, compared with $2{\varepsilon_{0}}/E_{\rm F}$ (dotted line). (b): Composite-fermion number density ${n_{\mathrm{CF}}}$ normalized to ${n_{\mathrm{F}}}$ (full lines), as obtained from Eq. (III.1) for different values of the bosonic concentration ($x=1.0,0.76,0.5,0.175$, from top to bottom) compared with ${C_{\mathrm{BF}}}/(8\pi m_{r}{\varepsilon_{0}}{n_{\mathrm{F}}})$ (dotted lines) and $({n_{\mathrm{B}}}-n_{0})/{n_{\mathrm{F}}}$ (dashed lines).
We show the results for ${C_{\mathrm{BF}}}$ of Eq. (62) in Fig. 15(a) from weak to strong BF attraction for different values of the bosonic concentration. The same quantity once divided by the boson density, ${C_{\mathrm{BF}}}/{n_{\mathrm{B}}}$, displays a universal behavior (see inset), in line with what already discussed for the condensate fraction (sec. IV.4) and bosonic momentum distribution (sec. IV.2). For this quantity, one sees that universality occurs for all couplings.
In the strong coupling regime, ${C_{\mathrm{BF}}}$ is seen in Fig. 15(a) to diverge exponentially with coupling. Specifically, as shown in the inset, the dimensionless quantity ${C_{\mathrm{BF}}}/(m_{\rm F}^{2}E_{\rm F}^{2})$ scales like $2{\varepsilon_{0}}/E_{\rm F}$ (dotted line in the inset), with the binding energy ${\varepsilon_{0}}/{E_{\mathrm{F}}}=2\mathrm{e}^{2g}$. In this limit, once $T({\mathbf{P}},\Omega)$ is replaced in Eq. (62) by $T_{\mathrm{SC}}({\mathbf{P}},\Omega)$, one has indeed
${C_{\mathrm{BF}}}$ $\approx 4m_{r}^{2}\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\int\frac{d\Omega}{2\pi}T_{\mathrm{SC}}({\mathbf{P}},\Omega)e^{i\Omega 0^{+}}$
$=8\pi m_{r}{\varepsilon_{0}}\int\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\int\frac{d\Omega}{2\pi}G_{\mathrm{CF}}({\mathbf{P}},\Omega)e^{i\Omega 0^{+}}$
@@
thus showing that in the strong coupling limit ${C_{\mathrm{BF}}}$ becomes proportional to the binding energy ${\varepsilon_{0}}$ multiplied by the density of BF pairs $n_{\mathrm{CF}}$ (where $n_{\mathrm{CF}}$ is defined by the strong-coupling expression (III.1)).
The presence of $n_{\mathrm{CF}}$ in Eq. (IV.5) suggests comparing $n_{\mathrm{CF}}$ as defined by Eq. (III.1) and by expression ${C_{\mathrm{BF}}}/(8\pi m_{r}{\varepsilon_{0}}{n_{\mathrm{F}}})$, in order to see when deviations from the asymptotic expression occur. This is done in Fig. 15(b) for several values of the concentration $x$ (full lines and dotted lines, respectively). One sees that deviations are most significant at small concentrations, but for $g\gtrsim 2.5$ the contact essentially coincides with its asymptotic expressions for all cases. We notice furthermore that for $g\gtrsim 1$, the quantity $n_{\mathrm{CF}}$ essentially coincides with the number of bosons out of the condensate, $n_{\mathrm{B}}-n_{0}$, (represented by the dashed lines in Fig. 15(b)). This justifies the assumption ${n_{\mathrm{CF}}}\simeq{n_{\mathrm{B}}}-n_{0}$ which we made in Sec. III.4 when discussing the strong-coupling limit of our theory.
IV.6 Identification of different physical regions in the coupling-concentration plane
Based on the results obtained in Sec. III and Sec. IV.2, the coupling ($g=-\ln(k_{\rm F}a_{\rm BF})$) vs. boson concentration ($x=n_{\rm B}/n_{\rm F}$) plane of a BF mixture can be divided in different physical region.

### v2

@@ 4.2 Boson and fermion momentum distributions
For generic coupling strength, the numerical Fermi momentum distributions are shown in Fig. 12 at different bosonic concentrations $x$. We only report the case $\eta_{\mathrm{B}}=0$, since the effect of the BB repulsion is only limited to a slight renormalization of the Fermi step. We find a qualitatively analogous phenomenology to Fig. 11 even though the latter is obtained in the strong-coupling limit of the theory. This is because the effective interaction acting on fermions is given by an interplay between $g$ and $x$, that is, even if $g$ is strong, reducing $x$ results in an effectively weaker $g$. In addition, we observe that the position of the Fermi step does not change significantly from weak to strong-coupling, in contrast to the 3D case where the Fermi sphere is emptied by the vanishing of the Fermi wave vector [86]. In contrast, in 2D the destruction of the Fermi surface occurring with coupling or concentration takes place due to a strong renormalization of the fermionic quasi-particles, as described in Secs. 3.2 and 3.4, and it is their weight rather than their associated Fermi momentum that vanishes.
@@
4.3 Boson and fermion chemical potentials
@@
Figure 13: (a): Bosonic chemical potential $\mu_{\mathrm{B}}$ (shifted by the binding energy ${\varepsilon_{0}}$), in units of $E_{\mathrm{F}}$, as a function of the BF coupling $g$, for different values of bosonic concentration $x$, and BB repulsion $\eta_{\mathrm{B}}=0$. Dotted lines on the left: weak-coupling expression (54). Dotted lines on the right: strong-coupling approximation of Sec. 3. Inset: comparison between $-\mu_{\mathrm{B}}$ and the binding energy ${\varepsilon_{0}}$ (dotted line), both in units of ${E_{\mathrm{F}}}$, as a function of $g$. (b): Corresponding fermionic chemical potential $\mu_{\mathrm{F}}$ in units of $E_{\mathrm{F}}$. Dotted lines on the left: weak-coupling expression (55) (reported for clarity only for the two extreme values of $x$). Dotted lines on the right: strong-coupling approximation of Sec. 3.
@@
We stress that the above perturbative expressions have been tested against independent fixed-node diffusion Quantum Monte Carlo calculations in [104]. The first term in Eq. (54) arises from the standard Bogoliubov theory in 2D [94] whilst the second contribution stems from the BF interaction. One sees in Fig. 13 that in the weak-coupling regime, the numerical curves for both chemical potentials neatly approach the asymptotic limits (54,55) (dotted line on the left side). For the bosonic chemical potential, this occurs in a universal way, i.e., independently of the value of $x$, as required by Eq. (54) which does not depend on the bosonic concentration (and ${\varepsilon_{0}}$ is exponentially suppressed in this regime). In the opposite regime, both chemical potentials approach their strong-coupling benchmarks, depicted by the dotted lines on the right side of the graph, which are seen to be strongly dependent on concentration. Note that, for the boson chemical potential, Fig. 13(a) reports the quantity ${\mu_{\mathrm{B}}}+{\varepsilon_{0}}$ in order to remove from the bosonic chemical potential the leading strong-coupling contribution (given by $-{\varepsilon_{0}}$). The inset avoids this subtraction and illustrates on a logarithmic scale how the bosonic chemical potential approaches the binding energy as $g$ increases, thus showing that all bosons are paired with fermions at large values of $g$.
[…]
@@
Finally, we note that we have verified that at finite $\eta_{\mathrm{B}}=0.1$ the results remain qualitatively the same, with quantitative changes affecting only ${\mu_{\mathrm{B}}}$, as also expected from the weak-coupling expression (54). The fermionic chemical potential is instead unaffected, in line with Eq. (55). On the other hand, in the strong-coupling regime, the boson-boson repulsion is essentially irrelevant due to the dominant dimer-atom repulsion. In summary, our results for the chemical potentials show that the present approach is able to recover both the pairing of bosons with fermions into composite fermions in the strong-coupling limit and the perturbative results of Ref. [104] in the opposite weak-coupling limit.
@@
Figure 14: (a): Bosonic condensate fraction $n_{0}/n_{\mathrm{B}}$ as a function of the BF coupling parameter $g$, for different values of boson concentration $x$ and BB repulsion $\eta_{\mathrm{B}}=0$. Circles: Diagrammatic Monte Carlo results for the polaron quasiparticle residue $Z$ [99]. Squares: non-self-consistent $T$-matrix approximation results for the polaron quasiparticle residue $Z$ [121]. Diamonds: non-self-consistent $T$-matrix approximation results for the polaron quasiparticle residue $Z$ by us. (b): Square of hybridization energy scale $\Delta_{0}^{2}=2\pi{\varepsilon_{0}}n_{0}/m_{r}$ (in units of $E_{\rm F}^{2}$), as a function of the BF coupling $g$, for different values of the bosonic concentration $x$ and boson-boson repulsion $\eta_{\mathrm{B}}=0$. Strong-coupling benchmarks from Sec. 3 are also reported as dotted lines on the right.
@@
4.4 Boson condensate density and atom-molecule hybridization energy scale
@@
Figure 14(a) also reports the results for the polaron quasiparticle residue $Z$ previously obtained with Diagrammatic Monte Carlo simulations (circles) [99] and with the non-self-consistent $T$-matrix approach [121] (squares). This is because in 3D it was found [91, 33] that the universal condensate fraction essentially matches the quasiparticle residue $Z$ of the Fermi polaron (which in our Bose-Fermi mixture corresponds to the limit of a single boson immersed in a Fermi gas).
@@
Such identification between the universal condensate fraction and $Z$ is challenged by the present results in 2D. One sees that, even though the condensate fraction $n_{0}/n_{\rm B}$ and the polaron quasiparticle residue $Z$ share the same qualitative behavior, significant quantitative discrepancies between the two quantities are present. We emphasize that the existence of such a discrepancy cannot be ascribed to the specific choice of the self-energy used in our calculation. One sees indeed that such a discrepancy occurs also when $Z$ is obtained with the non-self-consistent $T$-matrix approach (squares in Fig. 14(a)), which is exactly the same approach used here when carried over to the single-impurity limit. As an independent check, we have also calculated $Z$ from the derivative of the bosonic self-energy with respect to the frequency by using the same code used at finite $x$, adapted to the single-impurity limit. Our results (diamonds) fully agree with the corresponding results for $Z$ from [121], thus showing that the discrepancy between $Z$ and the condensate fraction cannot be ascribed to possible numerical errors. We thus conclude that the equivalence between the quasiparticle residue $Z$ of the Fermi polaron and the universal condensate fraction observed in 3D is only approximate, and the differences between these two quantities are amplified in 2D.
As a matter of fact, and contrary to what it was argued in [91], there is no reason why the limit for $x\to 0$ of the condensate fraction and the polaron residue $Z$ should coincide. The first quantity is defined by the ratio between the number of condensed bosons $N_{0}$ and the total number of bosons $N_{\rm B}$ in the thermodynamic limit ($N_{\rm B}\to\infty$, $V\to\infty$, at fixed $n_{\rm B}$ and $x$, $V$ being the volume). It is only in this limit that a condensate fraction is well defined. So, no matter how small $x$ is, the number of bosons $N_{\rm B}$ will always scale to infinity in the thermodynamic limit, while in the polaron limit one has instead $N_{\rm B}=1$. Therefore, the occupancy of the zero-momentum state for a single impurity cannot be in general related to the ratio $N_{0}/N_{\rm B}$ in the limit $x\to 0$. It is a matter of order of limits. In the first case (condensate fraction), one first fixes $x$, takes the thermodynamic limit $V\to\infty$, and finally let $x\to 0$. In the second case (polaron $Z$), one first fixes $N_{\rm B}=1$ and then takes the limit $V\to\infty$ in such a way that $x=1/(n_{\rm F}V)\to 0$. It is thus clear that in this second case the thermodynamic limit and the limit $x\to 0$ are taken simultaneously, while in the first case the thermodynamic limit is taken first, followed by the limit $x\to 0$.
A further difference with respect to the 3D case is observed in the behavior of $n_{0}$ for strong coupling. While in 3D $n_{0}$ drops identically to zero for values of $g$ larger than a critical coupling $g_{c}$, in 2D $n_{0}$ never vanishes (even though it becomes exponentially small with the coupling $g$). This in turn implies that the hybridization energy scale $\Delta_{0}\propto(n_{0}{\varepsilon_{0}})^{1/2}$ remains in general finite even at large values of $g$, except for the case $x=1$ (besides the trivial case $x=0$), as shown in Fig. 14(b). The absence of a critical coupling in 2D here obtained recalls the absence of the polaron-to-molecule transition which, as mentioned in Sec. 2.2, is a known shortcoming of the present approach when applied to the single impurity problem in 2D [97, 98]. This defect originates from an overestimate of the atom-molecule repulsion in the strong-coupling limit of the theory [97]. Such an overestimate persists also at finite $x$ (more on this point in Sec. 5). We cannot therefore exclude that the absence of a critical coupling strength in 2D is just an artifact of the present approximation, in particular at small values of $x$, which are closer to the single impurity limit.
4.5 Tan’s contact parameter and composite-fermion density
@@
For large momenta, the fermionic and bosonic distribution functions have the following asymptotic behaviors
$n_{\rm{F}}^{k\gg k_{\mathrm{F}}}\left(k\right)\simeq\frac{{{C_{\mathrm{BF}}}/4m_{r}^{2}}}{\left(\frac{k^{2}}{2m_{r}}-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}\right)^{2}},$ (56)
$n_{\rm{B}}^{k\gg k_{\mathrm{F}}}\left(k\right)\simeq\frac{{{C_{\mathrm{BF}}}/4m_{r}^{2}}}{\left(\frac{k^{2}}{2m_{r}}-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}\right)^{2}}+\frac{\Sigma^{2}_{12}}{4\left(\frac{k^{2}}{2m_{\mathrm{B}}}-{\mu_{\mathrm{B}}}\right)^{2}},$ (57)
@@
whose leading order term, proportional to $k^{-4}$, defines the Tan’s contact parameter $C_{s}=n_{s}(k)k^{4}$, $s=\mathrm{F,B}$, for fermions and bosons respectively [116, 117, 118]. The coefficient ${C_{\mathrm{BF}}}$ gives the contribution to Tan’s contacts due to BF pairing. Within our $T$-matrix self-energy approach, it can be shown to be given by
${C_{\mathrm{BF}}}$ $=4m_{r}^{2}\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\int\!\!\!\frac{d\Omega}{2\pi}T({\mathbf{P}},\Omega)e^{i\Omega 0^{+}},$ (58)
@@
Figure 15: (a): Dimensionless Tan’s contact constant ${C_{\mathrm{BF}}}/m_{\mathrm{F}}^{2}E_{\mathrm{F}}^{2}$ as a function of the BF coupling $g$, for different values of the bosonic concentration $x$ and BB repulsion $\eta_{\mathrm{B}}=0$. Inset: same data divided by the boson concentration $x$, compared with $2{\varepsilon_{0}}/E_{\rm F}$ (dotted line). (b): Composite-fermion number density ${n_{\mathrm{CF}}}$ in units of ${n_{\mathrm{F}}}$ (full lines), as obtained from Eq. (33) for different values of the bosonic concentration ($x=1.0,0.76,0.5,0.175$, from top to bottom) compared with ${C_{\mathrm{BF}}}/(8\pi m_{r}{\varepsilon_{0}}{n_{\mathrm{F}}})$ (dotted lines) and $({n_{\mathrm{B}}}-n_{0})/{n_{\mathrm{F}}}$ (dashed lines).
We show the results for ${C_{\mathrm{BF}}}$ of Eq. (58) in Fig. 15(a) from weak to strong BF attraction for different values of the bosonic concentration. The same quantity once divided by the boson density, ${C_{\mathrm{BF}}}/{n_{\mathrm{B}}}$, displays a universal behavior (see inset), in line with what already discussed for the condensate fraction (sec. 4.4) and bosonic momentum distribution (sec. 4.2). For this quantity, one sees that universality occurs for all couplings.
In the strong-coupling regime, ${C_{\mathrm{BF}}}$ is seen in Fig. 15(a) to diverge exponentially with coupling. Specifically, as shown in the inset, the dimensionless quantity ${C_{\mathrm{BF}}}/(m_{\rm F}^{2}E_{\rm F}^{2})$ scales like $2{\varepsilon_{0}}/E_{\rm F}$ (dotted line in the inset), with the binding energy ${\varepsilon_{0}}/{E_{\mathrm{F}}}=2\mathrm{e}^{2g}$. In this limit, once $T({\mathbf{P}},\Omega)$ is replaced in Eq.  (58) by $T_{\mathrm{SC}}({\mathbf{P}},\Omega)$, one has indeed
${C_{\mathrm{BF}}}$ $\approx 4m_{r}^{2}\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\int\!\!\!\frac{d\Omega}{2\pi}T_{\mathrm{SC}}({\mathbf{P}},\Omega)e^{i\Omega 0^{+}}$ (59)
$=8\pi m_{r}{\varepsilon_{0}}\int\!\!\!\frac{d{\mathbf{P}}}{(2\pi)^{2}}\int\!\!\!\frac{d\Omega}{2\pi}G_{\mathrm{CF}}({\mathbf{P}},\Omega)e^{i\Omega 0^{+}}$ (60)
@@
thus showing that in the strong-coupling limit ${C_{\mathrm{BF}}}$ becomes proportional to the binding energy ${\varepsilon_{0}}$ multiplied by the density of BF pairs $n_{\mathrm{CF}}$ (where $n_{\mathrm{CF}}$ is defined by the strong-coupling expression (33)).
The presence of $n_{\mathrm{CF}}$ in Eq. (59) suggests comparing $n_{\mathrm{CF}}$ defined by Eq. (33) to the expression ${C_{\mathrm{BF}}}/(8\pi m_{r}{\varepsilon_{0}}{n_{\mathrm{F}}})$, in order to see when deviations from the asymptotic expression occur. This is done in Fig. 15(b) for several values of the concentration $x$ (full lines and dotted lines, respectively). One sees that deviations are most significant at small concentrations, but for $g\gtrsim 2.5$ the contact essentially coincides with its asymptotic expression for all cases. We notice furthermore that for $g\gtrsim 1$, the quantity $n_{\mathrm{CF}}$ essentially coincides with the number of bosons out of the condensate, $n_{\mathrm{B}}-n_{0}$, (represented by the dashed lines in Fig. 15(b)). This justifies the assumption ${n_{\mathrm{CF}}}\simeq{n_{\mathrm{B}}}-n_{0}$ which we made in Sec. 3.4 when discussing the strong-coupling limit of our theory.
4.6 Identification of different physical regions in the coupling-concentration plane
Based on the results obtained in Secs. 3.23.44.3, and 4.4, the coupling ($g=-\ln(k_{\rm F}a_{\rm BF})$) vs. boson concentration ($x=n_{\rm B}/n_{\rm F}$) plane of a BF mixture can be divided in different physical region.